# AMUN: Adversarial Machine UNlearning

### v2

@@ Appendix F Comparison Using Prior Evaluation Methods
In § 6.3, we showed AMUN, whether with adaptive computation of $\mathcal{D_{\text{A}}}$ or using the pre-computed ones, outperforms other unlearning methods when handling multiple unlearning requests. Another important observation on the presented results in Figure 2 is that the effectiveness of unlearning decreases with the number of unlearning requests. For the setting with access to $\mathcal{D_{\text{R}}}$, this decrease is due to the fact that the $\mathcal{D_{\text{F}}}$ at each step has been a part of $\mathcal{D_{\text{R}}}$ at the previous steps; the model has been fine-tuned on this data in all the previous steps which has led to further improving confidence of the modes on predicting those samples. This result also matches the theoretical and experimental results in differential privacy literature as well [Dwork, 2006, Abadi et al., 2016].
This problem does not exist for the setting where there is no access to $\mathcal{D_{\text{R}}}$, but we still see a decrease in the unlearning effectiveness as we increase the number of unlearning requests. The reason behind this deterioration is that the model itself is becoming weaker. As Figure 6 shows, the accuracy on the model on both $\mathcal{D_{\text{R}}}$ and $\mathcal{D_{\text{T}}}$ gets worse as it proceeds with the unlearning request; this is because each unlearning step shows the model only $2\%$ ($1$K) of the samples and their corresponding adversarial examples for fine-tuning. So this deterioration is expected after a few unlearning requests. So when using AMUN in this setting (no access to $\mathcal{D_{\text{R}}}$) in practice, it would be better to decrease the number of times that the unlearning request is performed, for example by performing a lazy-unlearning (waiting for a certain number of requests to accumulate) or at least using a sub-sample of $\mathcal{D_{\text{R}}}$ if that is an option.

### v3

@@ Appendix F Comparison Using Prior Evaluation Methods
In § 6.3, we showed AMUN, whether with adaptive computation of $\mathcal{D_{\text{A}}}$ or using the pre-computed ones, outperforms other unlearning methods when handling multiple unlearning requests. Another important observation on the presented results in Figure 2 is that the effectiveness of unlearning decreases with the number of unlearning requests. For the setting with access to $\mathcal{D_{\text{R}}}$, this decrease is due to the fact that the $\mathcal{D_{\text{F}}}$ at each step has been a part of $\mathcal{D_{\text{R}}}$ at the previous steps; the model has been fine-tuned on this data in all the previous steps which has led to further improving confidence of the modes on predicting those samples. This result also matches the theoretical and experimental results in differential privacy literature as well [dwork2006differential, abadi2016deep].
This problem does not exist for the setting where there is no access to $\mathcal{D_{\text{R}}}$, but we still see a decrease in the unlearning effectiveness as we increase the number of unlearning requests. The reason behind this deterioration is that the model itself is becoming weaker. As Figure 6 shows, the accuracy on the model on both $\mathcal{D_{\text{R}}}$ and $\mathcal{D_{\text{T}}}$ gets worse as it proceeds with the unlearning request; this is because each unlearning step shows the model only $2\%$ (11K) of the samples and their corresponding adversarial examples for fine-tuning. So this deterioration is expected after a few unlearning requests. So when using AMUN in this setting (no access to $\mathcal{D_{\text{R}}}$) in practice, it would be better to decrease the number of times that the unlearning request is performed, for example by performing a lazy-unlearning (waiting for a certain number of requests to accumulate) or at least using a sub-sample of $\mathcal{D_{\text{R}}}$ if that is an option.